Calculate the pH of the solution after the addition of the following amounts of 0.0676 M HNO3 to a 60.0 mL solution of 0.0750 Maziridine. The pKa of aziridinium is 8.04 d) 62.6 mL of HNO3 a) 0.00 mL of HNO3 Number Number pH pH e) Volume of HNO3 equal to b) 6.81 mL of HNO3 the equivalence point Number Number pH pH c) Volume of HNO3 equal to half f) 70.5 mL of HNO3 the equivalence point volume Number Number pH pH
Answer #1

pH calculation

a) 0.00 ml HNO3 added

[aziridine] = 0.0750 M

aziridine + H2O <==> aziridineH+ + OH-

let x amount has reacted

Kb = [aziridineH+][OH-]/[aziridine]

pKa = 8.04

pKb = 14 - pKa = 5.96

pKb = -log[Kb] = 5.96

Kb = 1.1 x 10^-6

1.1 x 10^-6 = x^2/0.0750

x = [OH-] = 2.87 x 10^-4 M

pOH = -log[OH-] = 3.54

pH = 14 - pOH = 10.46

b) 6.81 ml HNO3 added

initial moles [aziridine] = 0.0750 M x 60 ml = 4.5 mmol

moles HNO3 added = 0.0676 M x 6.81 ml = 0.46 mmol

[aziridineH+] formed = 0.46 mmol/66.81 ml = 0.007 M [acid]

[aziridine] remained = (4.5 - 0.46) mmol/66.81 ml = 0.060 M [base]

pH = pKa + log(base/acid)

     = 8.04 + log(0.060/0.007)

     = 8.98

c) volume of HNO3 equal to half the equivalence point volume

initial moles [aziridine] = 0.0750 M x 60 ml = 4.5 mmol

Volume of HNO3 added = 4.5 mmol/2 x 0.0676 M = 33.28 ml

moles HNO3 added = 0.0676 M x 33.28 ml = 2.25 mmol

[aziridineH+] formed = 2.25 mmol/93.28 ml = 0.024 M [acid]

[aziridine] remained = (4.5 - 2.25) mmol/93.28 ml = 0.024 M [base]

pH = pKa + log(base/acid)

     = 8.04 + log(0.024/0.024)

     = 8.04

at this point half of aziridine has reacted with added HNO3 to form aziridineH+

So,

[aziridineH+] formed = [aziridine] remained in solution

pH = pKa = 8.04

d) 62.6 ml HNO3 added

initial moles [aziridine] = 0.0750 M x 60 ml = 4.5 mmol

moles HNO3 added = 0.0676 M x 62.6 ml = 4.232 mmol

[aziridineH+] formed = 4.232 mmol/122.6 ml = 0.0345 M [acid]

[aziridine] remained = (4.5 - 4.232) mmol/122.6 ml = 0.0022 M [base]

pH = pKa + log(base/acid)

     = 8.04 + log(0.0022/0.0345)

     = 6.84

e) Volume of HNO3 equal to the equivalence point

initial moles [aziridine] = 0.0750 M x 60 ml = 4.5 mmol

Volume of HNO3 added to reach equivalence point = 4.5 mmol/0.0676 M = 66.57 ml

moles HNO3 added = 0.0676 M x 33.28 ml = 4.5 mmol

all of aziridine is neutralized

[aziridineH+] formed = 4.5 mmol/126.57 ml = 0.0355 M [acid]

aziridineH+ + H2O <==> aziridine + H3O+

let x amount has reacted

Ka = [aziridine][H3O+]/[aziridineH+]

pKa = -log[Ka] = 8.04

9.12 x 10^-9 = x^2/0.0355

x = [H3O+] = 1.80 x 10^-5 M

pH = -log[H3O+] = 4.745

f) 70.5 ml HNO3 added

initial moles [aziridine] = 0.0750 M x 60 ml = 4.5 mmol

moles HNO3 added = 0.0676 M x 70.5 ml = 4.766 mmol

excess [HNO3] = [H+] present

                         = (4.766 - 4.5) mmol/130.5 ml

                         = 2.04 x 10^-3 M

pH = -log[H+] = 2.69